# Dimensions underlying the representational alignment of deep neural networks with humans

### v1

@@ 1.3 Factors underlying similarities and differences between humans and DNNs
Figure 5: Factors that determine the similarity between human and VGG-16 embedding dimensions. a, Representational similarity matrices reconstructed from the human and VGG-16 embedding. b, Pairwise correlations between human and VGG-16 embedding dimensions. c, Cumulative RSA analysis that shows the amount of variance explained in the human RSM as a function of the number of DNN dimensions. The black line shows the number of dimensions required to explain 95% of the variance. d-f, Intersection (red and blue regions) and differences (orange and green regions) between three highly correlating human and DNN dimensions. For this figure, we filtered the embedding by images from the public domain.
@@
Given the imperfect alignment of DNN and human dimensions, we explored the similarities and differences in the stimuli represented by these dimensions. For each dimension, we identified which images were most representative of both humans and the DNN. Crucially, to highlight the discrepancies between the two domains, we then identified which images exhibited strong dimension values for humans but weak values for the DNN, and vice versa (Fig. 5d-f). While the results indicated similar visual and semantic representations in the most representative images, they also exposed clear divergences in dimension meanings. For instance, in an animal-related dimension, humans consistently represented animals even for images where the DNN exhibited very low dimension values. Conversely, the DNN dimension strongly represented objects that were not animals, such as natural objects, cages, or mesh (Fig. 5d). Similarly, a string-related dimension maintained a string-like representation in humans but included other objects in the DNN that were not string-like, potentially reflecting features related to thin, curvy objects or specific image features (Fig. 5f).
1.4 Relevance of object dimensions for DNN and human categorization behavior
@@ 2 Discussion
A key challenge in understanding the similarities and differences in humans and artificial intelligence (AI) lies in establishing ways to make these two domains directly comparable. Overcoming this challenge would allow us to identify strategies for making AI more human-like (Geirhos et al., 2018) and for using AI as effective models of human perception and cognition. In this work, we propose a framework to identify interpretable factors that determine the similarities and differences between human and AI representations. In this framework, these factors can be identified by using the same experiment to probe behavior in humans and AI systems and applying the same computational strategy to the natural and artificial responses for inferring their respective interpretable embeddings. We successfully applied this approach to human similarity judgments and representations in a deep neural network (DNN) trained to classify natural images, thus allowing for a direct, meaningful comparison of the representations between both domains.
@@
Our results revealed that the DNN contained representations that appeared to be similar to those found in humans, ranging from visual (e.g., "white", "circular/round", "transparent") to semantic features (e.g., "food-related", "fire-related"). However, a direct comparison to humans showed largely different strategies for arriving at these representations. While human representations were dominated by semantic dimensions, the DNN exhibited a pronounced bias towards visual or mixed visual-semantic dimensions. In addition, a face-to-face comparison of seemingly aligned dimensions revealed that DNNs only approximated the semantic representations found in humans. These different strategies were also reflected in their behavior, where similar behavioral outcomes were based on different embedding dimensions. Thus, despite seemingly well aligned human and DNN representations at a global level, deriving dimensions underlying the representational similarities provided a more complete and more fine-grained picture of this alignment, revealing the nature of the representational strategies that humans and DNNs use (Sucholutsky et al., 2023; Cichy and Kaiser, 2019; Kanwisher et al., 2023).
@@
While previous approaches like RSA (Kornblith et al., 2019; Kriegeskorte et al., 2008) are particularly useful for comparing one or multiple representational spaces, they typically only provide a global quantitative measure of alignment and require explicit hypotheses and comparisons to other explicit models of representation to determine what it is about the representational space that drives human alignment. In contrast, other approaches have focused specifically on the interpretability of DNN representations (Zhou et al., 2018; Morcos et al., 2018; Erhan et al., 2009; Mahendran and Vedaldi, 2015; Nguyen et al., 2019; Bau et al., 2017, 2020) but either provide very specific local measures about DNN units or have limited direct comparability to human representations, given that the same interpretability methods can typically not be applied to understand human mental representations. Our framework combines the strengths of the comparability gained from RSA and existing interpretability methods to understand image processing in DNNs. We applied common interpretability methods to show that our work allows for detailed experimental testing and causal probing of DNN representations and behavior in response to diverse images. Yet only the direct comparison to human representations revealed the diverging representational strategies of humans and DNNs and thus limitations of the visualization techniques we used (Geirhos et al., 2023).
Our results are consistent with previous work indicating that DNNs make use of "shortcut" strategies that deviate from those used in humans (Geirhos et al., 2020; Hermann et al., 2023). Beyond the existing known biases, here we found a visual bias in DNNs that diverges from a more semantic bias in humans that underlies similarity judgments. This semantic bias is in line with the known ability of humans to abstract away from their immediate percept and their ability to structure the visual world based on their knowledge about the objects and their meaning (Palmeri and Gauthier, 2004). In contrast, even the highest layers in DNNs carry representations that continue to exhibit striking visual biases to solve the tasks they had been trained on, including image classification or relating images to text. While the identification of the visual bias is useful for understanding differences in representations and behavior, future work needs to determine the foundation of this bias and what changes in architectures or training are required to reduce it.
@@
The framework introduced in this work can be expanded in multiple ways. Future work could translate this approach to a comprehensive overview across arbitrary DNN architectures, training objectives, or training datasets (Muttenthaler et al., 2023a, b; Conwell et al., 2022), which could reveal specific strategies for increasing representational alignment (Peterson et al., 2018; Fel et al., 2022). Varying the behavioral task beyond an odd-one-out task could further establish the alignment and misalignment according to behaviors beyond categorization. Finally, the framework could be applied to other domains, including brain recordings or other types of stimuli. Together, this framework promises a more comprehensive understanding of the relationship between human and AI representations, providing the potential to identify better candidate models of human cognition and behavior and more human-aligned artificial cognitive systems.
@@ 3.1 Triplet odd-one-out task
In the triplet odd-one-out task, participants are presented with three objects and must select the one that doesn’t fit. We define a dataset $\mathcal{D}\coloneqq\left\{\left(\{i_{s},j_{s},k_{s}\},\{a_{s},b_{s}\}\right)\right\}_{s=1}^{n}$ where $n$ is the total number of triplets and $\{i_{s},j_{s},k_{s}\}$ is a set of three unique objects, with $\{a_{s},b_{s}\}$ being the pair among them determined as most similar. We used a dataset of human responses collected by Hebart et al. (2020) to learn an embedding of human object concepts. In addition, we simulated the triplet choices from a DNN. For the DNN, we simulated these choices by computing the dot product of the penultimate layer activation ${\bm{z}}_{i}\in\mathbb{R}_{+}$ after applying the ReLU function, where $S_{ij}={\bm{z}}_{i}^{\top}{\bm{z}}_{j}$. The most similar pair $\{a_{s},b_{s}\}$ was then identified by the largest dot-product:
@@ 3.3 Embedding reproducibility and selection
We assessed reproducibility across 32 model runs with different seeds using a split-half reliability test. We chose the split-half reliability test for its effectiveness in evaluating the consistency of our model’s performance across different subsets of data, ensuring robustness. We partitioned the objects into two disjoint sets using odd and even masks. For each model run and every dimension in an embedding, we identified the dimension that is most highly correlated among all other models by using the odd-mask. Using the even mask, we correlated this highest match with the corresponding dimension. This process generated a sampling distribution of Pearson’s $r$ coefficients for all model seeds. We subsequently Fisher $z$-transformed the Pearson’s $r$ sampling distribution. The average $z$-transformed reliability score for each model run was obtained by taking the mean of these $z$-scores. Inverting this average provides an average Pearson’s $r$ reliability score (see Supplementary Information D). For our final model and all subsequent analyses, we selected the embedding with the highest average reproducibility across all dimensions.
@@ 3.4 Labeling dimensions and construction of word clouds
For the DNN, we labeled dimensions using human judgments. This allowed us to capture a broad and nuanced understanding of each dimension’s characteristics. To collect human judgments, we asked 12 laboratory participants (6 male, 6 female; mean age, 29.08; s.d. 3.09; range 25-35) to label each DNN dimension. Participants were presented with a $5\times 6$ grid of images, with each row representing a decreasing percentile of importance for that specific dimension. The top row contained the most important images, and the following rows included images within the 8th, 16th, 24th, and 32nd percentiles. Participants were asked to provide up to five labels that they thought best described each dimension. Word clouds showing the provided object labels were weighted by the frequencies of occurrence, and the top 6 labels were visualized.
@@ 3.5 Dimension Ratings
Two independent experts rated the dimensions according to two questions. The first question asked whether the dimensions were primarily visual-perceptual, semantic-conceptual, a mix of both, or whether their nature was unclear. For the second question, they rated the dimensions according to whether they reflected a single concept, several concepts, or were not interpretable. Overall, both raters agreed agreed 81.86% of the time for question 1 and 90.00% of the time for question 2. Response ambiguity was resolved by a third rater (see Supplementary Information A, B). All raters were part of the laboratory but were blind to whether the dimensions were model- or human-generated.
@@ 3.6 Dimension value maximization
Gradient-based optimization. To refine these initial images, we performed gradient-based optimization in the latent space of StyleGAN XL. Our objective function $\mathcal{L}\text{AM}$ balances two goals: increasing the absolute value of the embedding for dimension $j$ and concentrating probability mass towards dimension $j$. Formally, we define $\mathcal{L}_{\text{AM}}$ as:
@@
3.7 Highlighting image features
@@ Acknowledgments
FPM, LM and MNH acknowledge support by a Max Planck Research Group grant of the Max Planck Society awarded to MNH. MNH acknowledges support by the ERC Starting Grant COREDIM (ERC-StG-2021-101039712) and the Hessian Ministry of Higher Education, Science, Research and Art (LOEWE Start Professorship and Excellence Program “The Adaptive Mind”). U.G. acknowledges support from the project Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 of the research program Gravitation which is (partly) financed by the Dutch Research Council (NWO). LM acknowledges support by the German Federal Ministry of Education and Research (BMBF) for the Berlin Institute for the Foundations of Learning and Data (BIFOLD) (01IS18037A) and for the grants BIFOLD22B and BIFOLD23B. This study used the high-performance from the Raven and Cobra Linux clusters at the Max Planck Computing & Data Facility (MPCDF), Garching, Germany https://www.mpcdf.mpg.de/services/supercomputing/.
@@
Conceptualization: FPM, LM, MNH; Funding acquisition: MNH; Software: FPM, LM; Supervision: MNH; Visualization: FPM, LM; Writing – original draft: FPM, LM; Writing – final manuscript: FPM, LM, UG, MNH.
Code availability
A PyTorch implementation of the model alongside all experiments presented in this paper is publicly available at https://github.com/florianmahner/object-dimensions/.
@@ Appendix A Dimension ratings and RSA across models
Figure S1: Dimension ratings and representational similarity across models. a, VGG-16 does not perform poorly when compared to other models, including Resnet50, DenseNet, CLIP, and BarlowTwins-Resnet50. b, The visual bias identified in VGG-16 is also evident across these other architectures, demonstrating consistent differences between human and DNN dimensions.
@@
We used VGG-16 due to its common use in computational cognitive neuroscience. To validate that VGG-16 is a suitable choice for the comparison to humans, we conducted RSA analyses with various other DNN models that differ in training diets, objective functions, and architecture. Using thingsvision [Muttenthaler and Hebart, 2021], we similarly extracted the penultimate features of these models and learned representational embeddings based on simulated triplet choices. Each embedding was then compared to the human-derived one using RSA. In Figure S1a we can see that VGG-16 does not perform poorly compared to other architectures, which suggests that it is a suitable choice for our analyses. Additionally, we assessed the visual bias in these architectures by having human raters categorize each dimension’s dominant visual property as visual, semantic, a mixture of both, or unclear. This reveals that the visual bias we find for VGG-16 also replicates across different DNNs (Fig. S1b)
Appendix B Human ratings of dimension interpretability

### v2

@@ 1.3 Factors driving human and DNN similarities and differences
Figure 5: Factors that determine the similarity between human and VGG-16 embedding dimensions. a, Representational similarity matrices reconstructed from the human and VGG-16 embedding. Each row represents an object, with rows sorted into 27 superordinate categories (e.g., animal, food, furniture) from (Hebart et al., 2019) to better highlight similarities and differences in representation. b, Pairwise correlations between human and VGG-16 embedding dimensions. c, Cumulative RSA analysis that shows the amount of variance explained in the human RSM as a function of the number of DNN dimensions. The black line shows the number of dimensions required to explain 95% of the variance. d-f, Intersection (red and blue regions) and differences (orange and green regions) between three highly correlating human and DNN dimensions. Pink circles denote the intersection of the red and blue regions, i.e., where the same image scores highly in both dimensions. For this figure, we filtered the embedding by images from the public domain.
@@
Given the imperfect alignment of DNN and human dimensions, we explored the similarities and differences in the stimuli represented by these dimensions. For each dimension, we identified which images were most representative of both humans and the DNN. Crucially, to highlight the discrepancies between the two domains, we then identified which images exhibited strong dimension values for humans but weak values for the DNN, and vice versa (Fig. 5d-f). While the results indicated similar visual and semantic representations in the most representative images, they also exposed clear divergences in dimension meanings. For instance, in an animal-related dimension, humans consistently represented animals even for images where the DNN exhibited very low dimension values. Conversely, the DNN dimension strongly represented objects that were not animals, such as natural objects, cages, or mesh (Fig. 5d). Similarly, a string-related dimension maintained a string-like representation in humans but included other objects in the DNN that were not string-like, potentially reflecting properties related to thin, curvy objects or specific image properties (Fig. 5f).
1.4 Relevance of object dimensions for categorization behavior
@@ 2 Discussion
A key challenge in understanding the similarities and differences in humans and artificial intelligence (AI) lies in establishing ways to make these two domains directly comparable. Overcoming this challenge would allow us to identify strategies to make AI more human-like (Geirhos et al., 2018) and for using AI as effective models of human perception and cognition. In this work, we propose a framework to identify interpretable factors that determine the similarities and differences between human and AI representations. In this framework, these factors can be identified by using the same experiment to probe behavior in humans and AI systems and applying the same computational strategy to the natural and artificial responses to infer their respective interpretable embeddings. We applied this approach to human similarity judgments and representations of DNNs trained on natural images with varying objectives, with a primary focus on an image classification model. This allowed for a direct, meaningful comparison of the representations underlying human similarity judgments with the representations of the image classification model.
@@
Our results revealed that the DNN contained representations that appeared to be similar to those found in humans, ranging from visual (e.g., "white", "circular/round", "transparent") to semantic properties (e.g., "food-related", "fire-related"). However, a direct comparison to humans showed largely different strategies for arriving at these representations. While human representations were dominated by semantic dimensions, the DNN exhibited a pronounced bias towards visual or mixed visual-semantic dimensions. In addition, a direct comparison of seemingly aligned dimensions revealed that DNNs only approximated the semantic representations found in humans. These different strategies were also reflected in their behavior, where similar behavioral outcomes were based on different embedding dimensions. Thus, despite seemingly well aligned human and DNN representations at a global level, deriving dimensions underlying the representational similarities provided a more complete and more fine-grained picture of this alignment, revealing the nature of the representational strategies that humans and DNNs use (Cichy and Kaiser, 2019; Sucholutsky et al., 2023; Kanwisher et al., 2023).
@@
While approaches like RSA (Kriegeskorte et al., 2008; Kornblith et al., 2019) are particularly useful for comparing one or multiple representational spaces, they typically provide only a summary statistic of the degree of alignment and require explicit hypotheses and model comparisons to determine what it is about the representational space that drives human alignment. In contrast, other approaches have focused specifically on the interpretability of DNN representations (Erhan et al., 2009; Zeiler and Fergus, 2014; Mahendran and Vedaldi, 2015; Bau et al., 2017; Morcos et al., 2018; Nguyen et al., 2019; Bau et al., 2020), but either provide very specific local measures about DNN units or have limited direct comparability to human representations, as the same interpretability methods can typically not be applied to understand human mental representations. Our framework combines the strengths of the comparability gained from RSA and existing interpretability methods to understand image representations in DNNs. We applied common interpretability methods to show that our approach allows for detailed experimental testing and causal probing of DNN representations and behavior across diverse images. Yet, only the direct comparison to human representations revealed the diverging representational strategies of humans and DNNs and thus limitations of the visualization techniques we used (Geirhos et al., 2023).
Our results are consistent with previous work indicating that DNNs make use of strategies that deviate from those used in humans (Geirhos et al., 2020; Hermann et al., 2023). Beyond previously discovered biases, here we found a visual bias in DNNs that diverges from a semantic bias in humans for similarity judgments. Notably, even the highest layers in DNNs retained strong visual biases for solving the tasks they had been trained on, including image classification or linking images with text, which can both be described as semantic tasks with different degrees of richness. This visual strategy may, of course, reflect how our visual system solves core object recognition (DiCarlo et al., 2012). Indeed, it is an open question to what extent human core object recognition relies on a similar visual bias (Jagadeesh and Gardner, 2022) and whether this bias is also found in anterior ventral-temporal cortex (Prince et al., 2024), which is known to be involved in high-level object processing (Kanwisher, 2010) However, even if humans used a primarily visual strategy for solving core object recognition, our findings would demonstrate a significant limitation of DNNs in capturing human mental representations as measured with similarity judgments, despite similar representational geometries (Mur et al., 2013).
Interestingly, CLIP, a more predictive model of human cortical visual processing (Conwell et al., 2022; Wang et al., 2023), retained a visual bias despite training on semantic image descriptions, showing that classification objective alone is not sufficient for explaining visual bias in DNNs. At the same time, the visual bias in CLIP was smaller (see Supplementary Fig. 1b), indicating that better models of high-level visual processing may also be models with a larger semantic bias and pointing towards potential strategies for improving their alignment with humans, which may involve multimodal pretraining or larger, more diverse datasets (Wang et al., 2023). Future work would benefit from a systematic comparison of different DNNs to identify what factors determine visual bias and their alignment with human brain and behavioral data.
While these results indicate that studying core dimensions of DNN representations can improve our understanding of the factors required to identify better models of human mental representations, it has also been demonstrated recently that aligning DNNs with human representations can improve DNN robustness and performance at out-of-distribution tasks (Muttenthaler et al., 2023b; Sundaram et al., 2024). This work highlights that identifying visual bias may be useful not only for understanding representational and behavioral differences between humans and DNNs, but also for guiding future work determining the gaps in human-AI alignment and identifying adjustments in architecture, training needed to reduce this bias (Sucholutsky et al., 2023). Further work is needed to clarify the role of task instructions in human-AI alignment across diverse tasks and instruction (Dwivedi and Roig, 2019).
@@
The framework introduced in this work can be expanded in several ways. Future work could use this approach to explore what factors make DNNs similar or different from one another. A comprehensive analysis of various DNN architectures, objectives, or datasets (Muttenthaler et al., 2023a, b; Conwell et al., 2022) could uncover the factors underlying representational alignment, and extension to other stimuli, tasks, and domains, including brain recordings. Together, this framework promises a more comprehensive understanding of the relationship between human and AI representations, providing the potential to identify better candidate models of human cognition and behavior and more human-aligned artificial cognitive systems.
@@ 3.1 Triplet odd-one-out task
In the triplet odd-one-out task, participants are presented with three objects and must choose the one that is least similar to the others, i.e. the odd-one-out. We define a dataset $\mathcal{D}\coloneqq\left\{\left(\{i_{s},j_{s},k_{s}\},\{a_{s},b_{s}\}\right)\right\}_{s=1}^{n}$ where $n$ is the total number of triplets and $\{i_{s},j_{s},k_{s}\}$ is a set of three unique objects, with $\{a_{s},b_{s}\}$ being the pair among them determined as most similar. We used a dataset of human responses collected by Hebart et al. (2020) to learn an embedding of human object concepts. In addition, we simulated the triplet choices from a DNN. For the DNN, we simulated these choices by computing the dot product of the penultimate layer activation ${\bm{z}}_{i}\in\mathbb{R}_{+}$ after applying the ReLU function, where $S_{ij}={\bm{z}}_{i}^{\top}{\bm{z}}_{j}$. The most similar pair $\{a_{s},b_{s}\}$ was then identified by the largest dot-product:
@@ 3.3 Embedding reproducibility and selection
We assessed reproducibility across 32 model runs with different seeds using a split-half reliability test. We chose the split-half reliability test for its effectiveness in evaluating the consistency of our model’s performance across different subsets of data, ensuring robustness. We partitioned the objects into two disjoint sets using odd and even masks. For each model run and every dimension in an embedding, we identified the dimension that is most highly correlated among all other models by using the odd-mask. Using the even mask, we correlated this highest match with the corresponding dimension. This process generated a sampling distribution of Pearson’s $r$ coefficients for all model seeds. We subsequently Fisher $z$-transformed the Pearson’s $r$ sampling distribution. The average $z$-transformed reliability score for each model run was obtained by taking the mean of these $z$-scores. Inverting this average provides an average Pearson’s $r$ reliability score (see Supplementary Information G). For our final model and all subsequent analyses, we selected the embedding with the highest average reproducibility across all dimensions.
@@ 3.4 Labeling dimensions and construction of word clouds
For the DNN, we labeled dimensions using human judgments. This allowed us to capture a broad and nuanced understanding of each dimension’s characteristics. To collect human judgments, we asked 12 laboratory participants (6 male, 6 female; mean age, 29.08; s.d. 3.09; range 25-35) to label each DNN dimension. Participants were presented with a $5\times 6$ grid of images, with each row representing a decreasing percentile of importance for that specific dimension. The top row contained the most important images, and the following rows included images within the 8th, 16th, 24th, and 32nd percentiles. Participants were asked to provide up to five labels that they thought best described each dimension. Word clouds showing the provided object labels were weighted by the frequencies of occurrence, and the top 6 labels were visualized. Due to computer crashes during data acquisition, three participants had incomplete data (32%, 80% and 93%).
@@ 3.5 Dimension Ratings
Two independent experts rated the dimensions according to two questions. The first question asked whether the dimensions were primarily visual-perceptual, semantic-conceptual, a mix of both, or whether their nature was unclear. For the second question, they rated the dimensions according to whether they reflected a single concept, several concepts, or were not interpretable. Overall, both raters agreed agreed 81.86% of the time for question 1 and 90.00% of the time for question 2. Response ambiguity was resolved by a third rater (see Supplementary Information A, B, C). All raters were part of the laboratory but were blind to whether the dimensions were model- or human-generated.
3.6 Convolutional Embeddings
We additionally learned embeddings from early (convolutional block 1), middle (convolutional block 3) and late (convolution block 5) convolutional layers of VGG-16. For this, we applied global average pooling to the spatial dimensions of the feature maps and then sampled triplets from the averaged 1d representations.
@@ 3.7 Dimension value maximization
Gradient-based optimization. To refine these initial images, we performed gradient-based optimization in the latent space of StyleGAN XL. Our objective function $\mathcal{L_{\text{AM}}}$ balances two goals: increasing the absolute value of the embedding for dimension $j$ and concentrating probability mass towards dimension $j$. Formally, we define $\mathcal{L}_{\text{AM}}$ as:
@@
3.8 Highlighting image properties
@@ 3.9 RSA analyses
Data Availability
The images used in this study are from the THINGS object concept and image database, available at the OSF repository (https://osf.io/jum2f). All result files pertaining to this study are made publicly available via a separate OSF repository (https://osf.io/nva43/).
Code availability
A Python implementation of all experiments presented in this paper is publicly available at https://github.com/florianmahner/object-dimensions/ (Mahner et al., 2025).
@@ Acknowledgments
FPM, LM and MNH acknowledge support by a Max Planck Research Group grant of the Max Planck Society awarded to MNH. MNH acknowledges support by the ERC Starting Grant COREDIM (ERC-StG-2021-101039712) and the Hessian Ministry of Higher Education, Science, Research and Art (LOEWE Start Professorship and Excellence Program “The Adaptive Mind”). U.G. acknowledges support from the project Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 of the research program Gravitation which is (partly) financed by the Dutch Research Council (NWO). LM acknowledges support by the German Federal Ministry of Education and Research (BMBF) for the Berlin Institute for the Foundations of Learning and Data (BIFOLD) (01IS18037A) and for the grants BIFOLD22B and BIFOLD23B. This study used the high-performance from the Raven and Cobra Linux clusters at the Max Planck Computing & Data Facility (MPCDF), Garching, Germany https://www.mpcdf.mpg.de/services/supercomputing/. The funders had no role in study design, data collection and analysis, decision to publish or preparation of the manuscript.
@@
Conceptualization: FPM, LM, MNH; Funding acquisition: MNH; Software: FPM, LM; Supervision: MNH; Visualization: FPM, LM; Writing – original draft: FPM, LM; MNH Writing – final manuscript: FPM, LM, UG, MNH.
Competing interests
There are no financial and non-financial competing interests.
@@ Appendix A Dimension ratings and RSA across models
Figure 1: Dimension ratings and representational similarity across models. a, VGG-16 does not perform poorly when compared to other models, including Resnet50, DenseNet, CLIP, and BarlowTwins-Resnet50. b, The visual bias identified in VGG-16 is also evident across these other architectures, demonstrating consistent differences between human and DNN dimensions.
@@
We used VGG-16 due to its common use in computational cognitive neuroscience. To validate that VGG-16 is a suitable choice for the comparison to humans, we conducted RSA analyses with various other DNN models that differ in training diets, objective functions, and architecture. Using thingsvision [Muttenthaler and Hebart, 2021], we similarly extracted the penultimate features of these models and learned representational embeddings based on simulated triplet choices. Each embedding was then compared to the human-derived one using RSA. In Supplementary Fig. 1a we can see that VGG-16 does not perform poorly compared to other architectures, which suggests that it is a suitable choice for our analyses. Additionally, we assessed the visual bias in these architectures by having human raters categorize each dimension’s dominant visual property as visual, semantic, a mixture of both, or unclear. This reveals that the visual bias we find for VGG-16 also replicates across different DNNs (Supplementary Fig. 1b).
Appendix B Dimension ratings and RSA across layers
Figure 2: Dimension ratings and representational similarity across different VGG-16 layers. We compared dimensions learned from features extracted from early, middle, late and penultimate layers of VGG-16. a, The embedding learned from the penultimate layer representations has the largest representational alignment to human behavior. b, The visual bias was strongest in early layers of VGG-16 and semantic information is added in later layers. The fraction of semantic dimensions compared to visual dimensions is largest in the penultimate embedding.
We demonstrate that the visual bias was present throughout the network with a gradient from early to late layers, with early representations showing the largest visual bias and penultimate representations the smallest. In addition, the penultimate embedding also exhibited the strongest human alignment (Supplementary Fig. 2).
Appendix C Human ratings of dimension interpretability